likeabossssssss (likeabossssssss):

Write an equation and then solve the problem. Vick’s Vitamins had start-up expenses of $85,000 for research and $125,000 for the first year’s inventory. So far, the sales are $45,000. What do the sales have to be for the rest of the year for the company to break even? A. $65,000 B. $130,000 C. $165,000 D. $255,000 a?

likeabossssssss (likeabossssssss):

@jabez177

likeabossssssss (likeabossssssss):

@AloneS.

OpenStudy (faiqraees):

Total earned = total invested Money required + Money earned = Research + Inventory x+45000=85000+125000 Solve it

likeabossssssss (likeabossssssss):

x=40k+45k=85k+125k = 210k @FaiqRaees

OpenStudy (faiqraees):

x+45k = 210k Solve it now

likeabossssssss (likeabossssssss):

165k + 45k = 210k

OpenStudy (faiqraees):

so what's the value of x?

likeabossssssss (likeabossssssss):

165k

likeabossssssss (likeabossssssss):

is x

OpenStudy (faiqraees):

good thats your answer

likeabossssssss (likeabossssssss):

thank u can u cheek a copel more ?

OpenStudy (faiqraees):

sure